### v2

# GPS-Synchronized Monitoring of Core-collapse Supernova Bursts with PandaX-4T via Coherent Elastic Neutrino Nuclear Scattering Binyu Pang

GPS-Synchronized Monitoring of Core-collapse Supernova Bursts with PandaX-4T via Coherent Elastic Neutrino Nuclear Scattering
Binyu Pang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Zihao Bo
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Wei Chen
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xun Chen
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yunhua Chen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Chen Cheng
School of Physics, Beihang University, Beijing 102206, China
Xiangyi Cui
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Manna Deng
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Yingjie Fan
Department of Physics, Yantai University, Yantai 264005, China
Deqing Fang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Xuanye Fu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Zhixing Gao
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Yujie Ge
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Lisheng Geng
School of Physics, Beihang University, Beijing 102206, China
Peng Huanwu Collaborative Center for Research and Education, Beihang University, Beijing 100191, China
International Research Center for Nuclei and Particles in the Cosmos & Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 100191, China
Southern Center for Nuclear-Science Theory (SCNT), Institute of Modern Physics, Chinese Academy of Sciences, Huizhou 516000, China
Karl Giboni
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xunan Guo
School of Physics, Beihang University, Beijing 102206, China
Xuyuan Guo
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Zichao Guo
School of Physics, Beihang University, Beijing 102206, China
Chencheng Han
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Ke Han
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Changda He
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinrong He
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Houqi Huang
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Junting Huang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yule Huang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Ruquan Hou
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xiangdong Ji
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Yonglin Ju
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xiaorun Lan
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Chenxiang Li
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jiafu Li
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Mingchuan Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Peiyuan Li
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Shuaijie Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Tao Li
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Yangdong Li
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Zhiyuan Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Qing Lin
State Key Laboratory of Particle Detection and Electronics, University of Science and Technology of China, Hefei 230026, China
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Jianglai Liu
jianglai.liu@sjtu.edu.cnState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yuanchun Liu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Congcong Lu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiaoying Lu
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Lingyin Luo
School of Physics, Peking University, Beijing 100871, China
Yunyang Luo
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Yugang Ma
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Yajun Mao
School of Physics, Peking University, Beijing 100871, China
Yue Meng
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Ningchun Qi
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Zhicheng Qian
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xiangxiang Ren
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Dong Shan
School of Physics, Nankai University, Tianjin 300071, China
Xiaofeng Shang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xiyuan Shao
School of Physics, Nankai University, Tianjin 300071, China
Guofang Shen
School of Physics, Beihang University, Beijing 102206, China
Manbin Shen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Wenliang Sun
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xuyan Sun
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Yi Tao
School of Science, Sun Yat-Sen University, Shenzhen 518107, China
Yueqiang Tian
School of Physics, Beihang University, Beijing 102206, China
Yuxin Tian
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Anqing Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Guanbo Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Hao Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Haoyu Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jiamin Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Lei Wang
College of Nuclear Technology and Automation Engineering, Chengdu University of Technology, Chengdu 610059, China
Meng Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Qiuhong Wang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Shaobo Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Shibo Wang
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Siguang Wang
School of Physics, Peking University, Beijing 100871, China
Wei Wang
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Xu Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Zhou Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yuehuan Wei
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Weihao Wu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yuan Wu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Mengjiao Xiao
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xiang Xiao
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Kaizhi Xiong
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Jianqin Xu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Yifan Xu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Shunyu Yao
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Binbin Yan
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Xiyu Yan
School of Physics and Astronomy, Sun Yat-Sen University, Zhuhai 519082, China
Yong Yang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Peihua Ye
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Chunxu Yu
School of Physics, Nankai University, Tianjin 300071, China
Ying Yuan
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Zhe Yuan
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Youhui Yun
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xinning Zeng
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Minzhen Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Peng Zhang
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Shibo Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Siyuan Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Shu Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Tao Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Wei Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Yang Zhang
yangzhangsdu@email.sdu.edu.cnResearch Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Yingxin Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Yuanyuan Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Li Zhao
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Kangkang Zhao
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Jifang Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Jiaxu Zhou
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Jiayi Zhou
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Ning Zhou
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xiaopeng Zhou
School of Physics, Beihang University, Beijing 102206, China
Zhizhen Zhou
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Chenhui Zhu
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
The landmark detection of neutrinos from SN1987A marked the dawn of neutrino astrophysics. The neutrino burst provided essential insights into fundamental properties of neutrinos, and served as key probes of stellar evolution and supernova dynamics. The recent advancement in coherent elastic neutrino-nucleus scattering enables the detection of core-collapse supernova burst neutrinos using tonne-scale liquid xenon detectors originally designed for dark matter direct detection. Leveraging this capability, we developed and deployed an online supernova monitoring system for the PandaX-4T experiment. This system features a GPS module with millisecond-level timing precision, a low false-alarm rate, and high sensitivity to galactic core-collapse supernova explosion events. The methodology is robust, directly scalable, and planned for implementation in the next-generation PandaX-20T experiment.
The core-collapse supernova (CCSN) explosions occur fewer than three times per century within the Milky Way [1]. The CCSN bursts occur in the latter stages of evolution of massive stars with masses exceeding eight times that of the Sun. Roughly erg of energy is released during a CCSN burst, with neutrinos carrying away about of the energy [2]. The neutrino burst lasts only a few or a few tens of seconds, and unlike electromagnetic radiation or gravitational waves, neutrinos can escape the stellar core more easily and are emitted earlier, allowing them to reach ground-based observatories before other messengers [3]. The only supernova (SN) neutrinos ever detected were from SN1987A, with around 20 events observed by a few detectors [4, 5, 6]. These neutrino events provided critical insights into the fundamental properties of neutrinos, and inspired theoretical models on CCSN explosion mechanisms [7, 8, 9]. However, progress in the field remains hindered by the lack of statistics of SN neutrinos. To prepare for neutrino detection from the next galactic CCSN burst, several neutrino experiments have implemented specialized supernova monitoring systems [10, 11, 12, 13], forming a global organization known as the Supernova Early Warning System (SNEWS) [14, 15].
Recently, significant advancements have been made in neutrino detection methods. In 2017, with neutrinos from a spallation neutron source, the COHERENT experiment reported the first observation of coherent elastic neutrino-nucleus scattering (CENS) [16], a coherently enhanced weak interaction predicted by Freedman in 1974 [17]. CENS offers a remarkable opportunity for detecting low energy astrophysical neutrinos [18, 19, 20, 21, 22], as its cross-section scales approximately with the square of the target nucleus’s neutron number, , where is the neutron number, is the Fermi coupling constant, and is the neutrino energy [23]. Furthermore, as a neutral current process, CENS is sensitive to all neutrino flavors, providing a valuable complement to experiments that primarily focus on specific flavors [24, 25]. The challenge, however, lies in the extremely low nuclear recoil energy, which is difficult to detect. Over the past two decades, liquid xenon-based dark matter direct detection experiments, designed to detect rare and low-energy scattering events, have made remarkable progress in improving their sensitivities. Experiments such as XENONnT, LUX-ZEPLIN (LZ), and XMASS [11, 26, 27] have proposed leveraging CENS to detect SN neutrinos. Notably, in 2024, the PandaX-4T [21] and XENONnT [22] collaborations independently reported indications of CENS from solar 8B neutrinos with xenon nuclei, and very recently, the LZ collaboration also reported the evidence of 8B CENS with more than 4- significance [28]. These findings mark a significant milestone towards the realization of SN neutrino detection via CENS.
In this paper, we will discuss the performance of a newly developed CCSN burst monitoring system in PandaX-4T,
utilizing online selection of CENS events.
The PandaX-4T experiment is located in the China Jinping Underground Laboratory (CJPL), shielded by 2,400 meters of rock overburden. Comprehensive details about the detector can be found in Ref. [29], with only a few essential aspects highlighted in this Letter. The primary detector is a cylindrical xenon time projection chamber (TPC) containing approximately 3.7 tonnes of liquid xenon (LXe). A physical event produces a delayed coincidence of photons from prompt scintillation (S1) and delayed electroluminescence photons from ionized electrons (S2), which are detected via the top and bottom photomultiplier (PMT) arrays. The waveforms from all fired PMTs are read out under a continuous and triggerless” data acquisition mode, and event building and reconstructions are performed by analysis software. The three-dimension interaction vertex is reconstructed using the top PMT pattern and the time difference between S1 and S2, allowing for further suppression of external background through fiducialization.
The PandaX-4T has been in operation since November 2020 and is scheduled to continue the data taking until the end of 2025. To allow an online monitoring for potential CCSN explosions, a dedicated GPS-based CCSN burst system was established in July 2025. About 83 days of physics data collected between July 18 and October 14, 2025, will be used to evaluate its performance.
A millisecond-level GPS-based timing system was specially designed for PandaX-4T, as shown in Fig. 1, to facilitate collaborative early warnings with other neutrino experiments via time-delay triangulation [15, 30]. The GPS ground sub-system comprises a GPS receiver and a camp control module. The GPS antenna receives satellite signals and transmits them to the GPS server, which generates a high-precision 10 MHz reference clock through its internal clocking mechanism to synchronize the local clock of the camp control module. The precise GPS time, encoded in the IRIG-B code and synchronized with the one pulse per second (PPS) signal, is transmitted to the camp control module and then downwards the control module in the underground laboratory (hereafter referred to as lab control module) through a multi-kilometer optical fiber network. A hardware clock counter of the lab control module synchronizes with the DAQ’s master clock at the start of each new data acquisition run (labeled as local clock and local run start in Fig. 1 ), and the mapping between the delivered GPS time and the hardware clock ticks is stored in a database of a time server. The time delay introduced by the fiber optic length is calculated by measuring the time difference between the emission and reception of synchronized signals sent in opposite directions (downward and upward time streams), using the camp and lab control modules. This delay is utilized to correct timestamps during data acquisition. The precision of this system is maintained at the millisecond level, which is determined primarily by the GPS server.
Fig. 1: Schematic diagram of the PandaX-4T’s GPS-based timing system, which comprises a GPS receiver, a camp control module, a lab control module and a time server. The mapping of the delivered GPS time and the DAQ hardware clock ticks is stored in a database. See text for more details. Fig. 2: Schematic diagram of SN online monitoring system in PandaX-4T. The time consumed in each step is shown.
We developed an online software-based monitoring system to identify CCSN bursts by detecting transient increases in CENS-like event rate in PandaX-4T induced by SN neutrinos. The flow diagram is shown in Fig. 2. When the data acquisition system generates a new raw data file, the acquired data is stored using the Bamboo-shoot3 library [31, 32]. Each raw file contains approximately one minute of data, with a size of about 2 GB. The data compression process reduces it to about 1 GB and writes the data to local disk, which takes about ten seconds.
The data is then transmitted to the remote data center in Chengdu, enabling parallel processing of different files. The transmission time is approximately fifteen seconds. Offline processing and event reconstruction take about 4.5 minutes. After reconstruction, event information such as S1 and S2 variables, reconstructed vertex, and other relevant parameters is extracted. A series of cuts are then applied to reduce background, including the afterglow cut (which removes periods immediately following large signals amounting to a 11% of deadtime), the fiducial volume (FV) cut, the accidental coincidence (AC) cut, and quality cuts for S1 and S2.
[…]
Fig. 3: The background event rate (a) and the false alert (b) as a function of time.
Finally, a CCSN burst alert is generated if the number of neutrino candidates exceeds the threshold () within a 10-second time window. Further details on this method are provided in Ref. [35]. The SN alert information, including the alert start time and the total number of neutrino candidates, is sent to the PandaX-4T SN group via email. This process takes approximately five seconds. Considering all latency factors discussed above, the entire monitoring process takes about 6 minutes, as also indicated in Fig. 2.
The operational status of the SN monitoring system is displayed on a webpage, allowing personnel to inspect key performance parameters. The system operates stably during routine data-taking periods, with exceptions occurring only during special circumstances, such as power outages or server maintenance. SN monitoring program also automatically excludes calibration data based on the data type defined by the data acquisition system.
To evaluate the efficiency of the single event selection described above, we calculate the expected neutrino energy spectrum using the so-called Garching model with progenitor masses of and ( is the solar mass), employing the LS220 nuclear equation of state (EoS) [36, 37]. The electroweak neutral current process allows most neutrinos to interact coherently with nucleons in the nucleus for neutrino energies MeV. Details of the differential cross-section calculation can be found in Ref. [35].
The signal response model for PandaX-4T is constructed using the Noble Element Simulation Technique (NEST) v2.3.6 parameterization [38], with relevant parameters tuned by fitting calibration data [39]. The NEST model simulates the light yield (, number of photons per keV) and charge yield (, number of electrons per keV) for nuclear recoil events in LXe. The S1 and S2 signals are then generated using a waveform simulation (WS) package, which assembles real waveforms from the data with and as inputs [40]. These WS events are reconstructed and selected using the same procedure as the real experimental data, allowing the overall detection efficiency to be determined.
The overall efficiency includes contributions from signal reconstruction, data quality selection, and the ROI cut, evaluated with the methodology outlined in Refs. [35, 41]. As an example, the overall detection efficiency for SN neutrino events using the Garching model with is approximately 13%. The most significant efficiency loss arises from the fact that the majority of CENS events have the low S1 and S2 values.
Based on the threshold mentioned above, eight alert candidates were observed during the 83 days of physics data. The time distribution of these alerts is shown in Fig. 3. Two alarms were identified in a same day, with a time difference of about 3 hours. No other anomalies were identified after a thorough check.
[…]
Based on this analysis, we conclude that within our 83-day dataset, the alerts are consistent with the absence of any real CCSN bursts above the background at the 1 level. It is important to note that the primary goal of SNEWS is to search for coincidences between multiple experiments, thereby significantly enhancing the sensitivity to CCSN signals beyond what is achievable by a single experiment. Notably, the false alert rate requirement for an experiment to join SNEWS is one per weeka criterion that PandaX-4T already satisfies.
[…]
Data type
Rate [/s]
Calendar time
Observed
Expected
DD
2.2
2.8 days
886
870.6
AmBe
9.7
3.7 days
288
265.3
Physical
2.9
83 days
8
6.1
For all eight alerts, intended alert messages to SNEWS were also generated, with one example alert shown in Table 2. More details about the SNEWS alert system can be found in Ref. [42].
For different (in total five) tiers of alerts, identified by the first column of Table 2, SNEWS provides different message template for participating experiments. For each tier, the second and third columns provide essential information for that example PandaX-4T alert.
Table 2: List of different message tiers within the SNEWS network, using a PandaX-4T alert as an example. The detector name is required in each tier and is not listed in the column.
For the Coincidence tier, the neutrino event time is represented in UTC format, derived from GPS time in PandaX-4T.
For the Significance tier, under the background only assumption, the p-values are the probabilities of observing more than or equal to the number of candidate events
within a 5-second time window at the single rate of .
For the Timing tier, the time series is an ordered sequence of nanosecond-precision integers, representing time offsets relative to the first neutrino candidate. For the Heartbeat tier, the detector status parameter is set to ON/OFF according to its operational condition.
For the Retraction tier, the retraction is initiated via a dedicated command provided by the SNEWS network, depending on real-time experimental considerations.
Tier
Required information
Value
Coincidence
initial neutrino time
2025-10-06T14:13:20.333181
Significance
p-values for time bins, width of time bins
[6.11e-06, 1.00], 5.00
Timing
initial neutrino time, neutrino time series
2025-10-06T14:13:20.333181, [0, 1213145668]
Heartbeats
detector status
ON/OFF
Retraction
retract latest
Based on the performance of the SN monitoring system described above, the detection potential of various CCSN explosions with PandaX-4T-only data can be evaluated. The differential event rate of CENS can be expressed as
where represents the time after the CCSN core bounce, tonnes is the effective target mass in this analysis, is Avogadro’s number, and the molar mass of xenon is , and is the distance from the CCSN to Earth. represents the minimum neutrino incident energy required to induce a nuclear recoil energy of in CENS. In practice, we integrate over MeV). and are the differential cross-section of the CENS process and the total neutrino fluxes of all flavors at the energy of and time of , respectively. represents the detection efficiency.
The expected number of SN neutrino events in the detector can be calculated by integrating Eq. 1. Here, we integrate the variable for the model with ( and ) over the time interval from the onset time of core bounce to 8.35 (7.60) seconds. is integrated up to 30 keV. The results are summarized in Table 3. The two used typical distances of 10 kpc and 168 pc correspond to the positions of the nearby Galactic center and Betelgeuse [43] as observed from Earth, respectively. The detection probability of CCSN explosions, defined as the fraction of the observed over the total, is estimated using the Garching model with , employing the LS220 EoS. Figure 4 illustrates the detection probability as a function of CCSN distance to the PandaX-4T and the next generation experiment PandaX-20T, where we assume that PandaX-20T has a fiducial mass of tonnes. PandaX-4T achieves nearly- detection efficiency at 10 kpc, while the PandaX-20T maintains 100% efficiency up to 30 kpc.
Table 3: The number of expected SN neutrinos for source distances of 10 kpc and 168 pc in PandaX-4T, using Garching models ( and ) with the LS220 EoS. The fourth column indicates the time interval from the onset time of core bounce to the model’s termination time.
SN model
d=10 kpc
d=168 pc
Duration [s]
11.2
3.7
1.3
7.60
27
8.1
2.9
8.35
Fig. 4: Detection probability of the CCSN explosions as a function of distance using the Garching model with for PandaX-4T (black) and PandaX-20T (red). The corresponding false alert rate is twice a month in PandaX-4T.
Fig. 5: Typical timing distributions of the detected SN neutrino events from the distance of 168 pc (a) and 10 kpc (b). Data points are produced by simulations and the red lines are the fit results. The fitted onset times relative to the true time in about 10,000 mock detections are shown in the bottom panels. The resulting onset time measurement exhibits a variation of ms (c) and s (d) relative to zero for 168 pc and 10 kpc, respectively, within a 68% confidence level (C.L.).
The core bounce onset time can also be determined when sufficient event statistics are available. A typical fitting result of the timings after the core bounce of the SN neutrino events from the distance of 168 pc is displayed in Fig. 5. The points represent the simulated SN neutrino events, and the histogram represents the fitting result using the equation , where denotes the time profile function of the CCSN events, while and are fit parameters.
To estimate how well we can determine the onset time, signified by the precision on the parameter , we repeat the fits for many copies of mock time profiles, with results shown in Figs. 5 and 5. The 68% C.L. intervals for the variation of the offset times relative to the true value are ms at 168 pc and s at 10 kpc due to smaller statistics. Note that the described fitting method may not represent the timing precision achievable for promptly sharing data from a real supernova, as the true burst profile may not match the input model. As an alternative, we also consider the core bounce onset time as the arrival time of the first neutrino candidate, based on sampling from the Garching model. The resulting 68% C.L. intervals for the onset time variation are approximately 1 ms at 168 pc and 0.2 s at 10 kpc. The uncertainty at 168 pc is dominated by GPS timing precision.
In summary, we have established a robust online supernova monitoring system with GPS time synchronization implemented in PandaX-4T, leveraging the novel detection channel of coherent elastic neutrino-nucleus scattering. The performance of this system during an 83-day data-taking period, from July 18, 2025, to October 14, 2025, is presented. During this period, eight alerts were produced, consistent with a false alert rate of 2.2 per month caused by background. This monitoring system has the capability to provide the astronomical community with early warnings of CCSN explosions with a latency of about 6 minutes, and achieves nearly detection efficiency for supernovae occurring within 10 kpc of Earth in PandaX-4T. The system is ready to join the SNEWS network.
Looking ahead, the implementation of this system in the next-generation PandaX-20T experiment is expected to significantly enhance sensitivity to core-collapse supernovae within our galaxy. By enabling precise measurements of the recoil energy spectrum and time profile features, it offers an exciting opportunity to significantly advance our understanding of the fundamental properties of neutrinos and provide critical insights for refining astrophysical models.
Acknowledgment: This project is supported in part by grants from National Key R&D Program of China (Nos. 2023YFA1606200, 2023YFA1606203), National Science Foundation of China (Nos. 12090060, 12090063, U23B2070, 12175139), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 21TQ1400218, 22JC1410100, 23JC1410200, ZJ2023-ZD-003). We thank for the support by the Discipline Construction Fund of Shandong University, and the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.